### v1

@@ Abstract
A $(\beta,\delta,\Delta)$-padded decomposition of an edge-weighted graph $G=(V,E,w)$ is a stochastic decomposition into clusters of diameter at most $\Delta$ such that for every vertex $v\in V$, the probability that $\mathbf{B}_{G}(v,\gamma\Delta)$ is entirely contained in the cluster containing $v$ is at least $e^{-\beta\gamma}$ for every $\gamma\in[0,\delta]$. Padded decompositions have been studied for decades and have found numerous applications, including metric embedding, multicommodity flow-cut gap, muticut, and zero extension problems, to name a few. In these applications, parameter $\beta$, called the padding parameter, is the most important parameter since it decides either the distortion or the approximation ratios. For general graphs with $n$ vertices, $\beta=\Theta(\log n)$.
@@ 1 Introduction
In this work, we study a stronger notion of stochastic decomposition, called padded composition. More formally, given a weighed graph $G=(V,E,w)$, a partition is $\Delta$-bounded if the diameter of every cluster is at most $\Delta$. A distribution $\mathcal{D}$ over partitions is called a $(\beta,\delta,\Delta)$-padded decomposition, if every partition is $\Delta$-bounded, and for every vertex $v\in V$ and $\gamma\in[0,\delta]$, we have:
$\Pr[\mathbf{B}_{G}(v,\gamma\Delta)\subseteq P(v)]\geq e^{-\beta\gamma}]\qquad\text{where $P(v)$ is the cluster containing $v$.}$ (1)
@@
A key new technical insight to all of our aforementioned results is the notion of tree-ordered net (Definition 2). A tree order net is analogous to the notion of nets, which were used extensively in designing algorithms for metric spaces. More formally, a $\Delta$-net is a set of points $N$ such that every two net points are at a distance at least $\Delta$ (i.e $\min_{x,y\in N}d_{X}(x,y)\geq\Delta$), and every point has a net point at a distance at most $\Delta$ (i.e. $\max_{x\in V}\min_{y\in N}d_{X}(x,y)\leq\Delta$). Filtser [Fil19a] showed that if there is a $\Delta$-net such that every ball of radius $3\Delta$ contains at most $\tau$ net points, that the metric admits a $\left(O(\log\tau),\Omega(1),O(\Delta)\right)$-padded decomposition. This result implies that metrics of doubling dimension $d$ have padding parameter $O(d)$ since doubling metrics have sparse nets: $\tau=2^{O(d)}$. Unfortunately, graphs of small treewidth do not have sparse nets; this holds even in very simple graphs such as star graphs. Nonetheless, we show that small treewidth graphs possess a structure almost as good: a net that is sparse w.r.t. some partial order. We formalize this property via tree-ordered nets. As we will later show, a sparse tree-ordered net is enough to construct the padded decomposition scheme; see Section 3. We believe that the notion of a tree-ordered net is of independent interest.
A tree order $\preceq$ of a set $V$ is a partial order (i.e. transitive, reflexive, and anti-symmetric) associated with a rooted tree $T$ and a map $\varphi:V\rightarrow V(T)$ such that $u\preceq v$ iff $\varphi(v)$ is an ancestor of $\varphi(u)$ in $T$. Given a weighted graph $G=(V,E,w)$, a tree order $\preceq$ w.r.t. tree $T$ is a valid order of $G$ if for every edge $\{u,v\}\in E$, it holds that $u\preceq v$ or $v\preceq u$ or both (i.e. $v$ is an ancestor of $u$, or vice versa). A simple consequence of the validity is that every connected subset $C$ in $G$ must contain a maximum element w.r.t $\preceq$; see 1. For a vertex $v\in V$, and subset $S\subseteq V$ let $S_{v\preceq}=\{u\in S\mid v\preceq u\}$ be the ancestors of $v$ w.r.t. $T$ in $S$. Similarly, let $S_{\preceq x}=\{u\in S\mid u\preceq x\}$ be all the descendants of $x$ in $S$. See the illustration on the right.
@@ 2 Preliminaries
We consider connected undirected graphs $G=(V,E)$ with edge weights $w:E\to\mathbb{R}_{\geq 0}$. We say that vertices $v,u$ are neighbors if $\{v,u\}\in E$. Let $d_{G}$ denote the shortest path metric in $G$. $\mathbf{B}_{G}(v,r)=\{u\in V\mid d_{G}(v,u)\leq r\}$ is the ball of radius $r$ around $v$. For a vertex $v\in V$ and a subset $A\subseteq V$, let $d_{G}(x,A):=\min_{a\in A}d_{G}(x,a)$, where $d_{G}(x,\emptyset)=\infty$. For a subset of vertices $A\subseteq V$, let $G[A]$ denote the induced graph on $A$, and let $G\setminus A:=G[V\setminus A]$.
@@
$\Pr[\mathbf{B}_{G}(z,\gamma\Delta)\subseteq P(z)]\geq e^{-\beta\gamma}\leavevmode\nobreak\ .$
@@ 3 Padded Decomposition and Sparse Cover
In this section, we prove three general lemmas constructing a padded decomposition and a sparse cover from a tree-ordered net. These lemmas, together with Lemma 1, imply Theorem 1, Theorem 2, and Theorem 3. We begin with Theorem 2, as its proof is simpler. We observe the following by the definition of a tree order.
Observation 1.
@@ 3.1 Proof of Theorem 2
$C_{i}=\mathbf{B}_{G[V_{\preceq x_{i}}]}(x_{i},\alpha\Delta)\leavevmode\nobreak\ ,$
@@
Finally, we show that for every vertex $v$, $\mathbf{B}_{G}(v,\frac{\Delta}{\beta})$ is fully contained in some cluster in $\mathcal{C}$, for $\beta=\frac{2}{\alpha-1}$. Let $B=\mathbf{B}_{G}(v,\frac{\Delta}{\beta})$. Let $v_{B}\in B$ be the closest vertex to the root w.r.t. $T$. 1 implies that for every vertex $u\in B$, $u\preceq v_{B}$. By the triangle inequality, for every $u\in B$ it holds that:
@@ 10             Add $C_{x}$ to $\mathcal{P}_{i}$
$d_{G}(x,x_{i})\leq d_{G[V_{\preceq x}]}(x,y_{i})+d_{G[V_{\preceq x_{i}}]}(y_{i},x_{i})\leq\alpha\Delta\leavevmode\nobreak\ .$
@@
$\leq\Delta+\frac{2\Delta}{\beta}=\left(1+\frac{2\cdot(\alpha-2)}{4}\right)\cdot\Delta=\frac{\alpha}{2}\cdot\Delta\leavevmode\nobreak\ ,$
@@ 3.3 Proof of Theorem 1
$C_{i}=\mathbf{B}_{G[V_{\preceq x_{i}}]}(x,R_{i})\setminus\cup_{j<i}C_{j}\leavevmode\nobreak\ .$
@@
Let $v_{B}\in B$ be the vertex closest to the root of $T$ (w.r.t distances in $T$). By 1, for every vertex $u\in B$, it holds that $u\preceq v_{B}$. Let $x_{B}\in N$ be the center that is an ancestor of $v_{B}$ and minimizes $\mathbf{B}_{G[V_{\preceq x}]}(x,v_{B})$. Note that $v_{B}$ will join the cluster of $x_{B}$, if it did not join any other cluster. It follows that no descendant of $x_{B}$ can be the center of the first cluster having a non-trivial intersection with $B$. This implies that for every center $x_{i}\notin N_{x_{B}\preceq}$, $\Pr[\mathcal{F}_{i}]=0$.
@@
We assume that by the round $i$, no vertex in $B$ is clustered, and that $d_{G[V_{\preceq x_{i}}]}(x_{i},B)\leq\beta\Delta$, as otherwise $\Pr[\mathcal{C}_{i}]=\Pr[\mathcal{F}_{i}]=0$ and we are done. Let $\rho$ be the minimal value of $\delta_{i}$ such that if $\delta_{i}\geq\rho$, some vertex of $B$ will join $C_{i}$. Formally $\rho=\frac{1}{\Delta}\cdot d_{G[V_{\preceq x_{i}}]}(x_{i},B)$. By our assumption, $\rho\leq\beta$. Set $\tilde{\rho}=\max\{\rho,1\}$. We have:
@@
$d_{G[V_{\preceq x_{i}}]}(u,x_{i})\leq d_{G[V_{\preceq x_{i}}]}(v_{i},x_{i})+2\gamma\Delta=\left(\rho+2\gamma\right)\cdot\Delta\leavevmode\nobreak\ .$
@@
$=\left(1-e^{-2\gamma\cdot\lambda}\right)\cdot\left(\Pr\left[\mathcal{F}_{i}\right]+\frac{1}{e^{(\beta-1)\cdot\lambda}-1}\right)\leavevmode\nobreak\ .$
@@
$\leq\left(1-e^{-2\gamma\cdot\lambda}\right)\cdot\left(1+e^{-2\gamma\cdot\lambda}\right)=1-e^{-4\gamma\cdot\lambda}\leavevmode\nobreak\ ,$
@@
$e^{-2\gamma\lambda}=\frac{e^{-2\gamma\lambda}\left(e^{(\beta-1)\cdot\lambda}-1\right)}{e^{(\beta-1)\cdot\lambda}-1}\geq\frac{e^{-2\gamma\lambda}\cdot e^{(\beta-1)\cdot\lambda}\cdot\frac{1}{2}}{e^{(\beta-1)\cdot\lambda}-1}\overset{(*)}{\geq}\frac{\frac{1}{2}\cdot e^{\frac{\alpha-1}{4}\cdot\lambda}}{e^{(\beta-1)\cdot\lambda}-1}=\frac{\tau}{e^{(\beta-1)\cdot\lambda}-1}\leavevmode\nobreak\ .$
@@ 4 Tree-Ordered Nets for Graphs of Bounded Tree-partition Width
The converse of 2 might not be true, specifically in the case when $u$ and $v$ are mapped to the same vertex in $\hat{T}$ by $\hat{\varphi}$. By 2, for any vertex $x\in V$, $V_{\preceq_{T}x}\subseteq V_{\preceq_{\hat{T}}x}$. Thus, for any $v$ and $x$ such that $v\preceq_{T}x$, if $d_{G}[V_{\preceq_{T}x}](v,x)\leq 3\Delta$, then $d_{G}[V_{\preceq_{\hat{T}}x}](v,x)\leq 3\Delta$. Therefore, $N^{3\Delta}_{v\preceq_{T}}\subseteq N^{3\Delta}_{v\preceq_{\hat{T}}}$ and hence $|N^{3\Delta}_{v\preceq_{T}}|\leq|N^{3\Delta}_{v\preceq_{\hat{T}}}|=\mathrm{poly}({\rm tp})$, giving the packing property.
Next, we show the covering property. By the covering property of $(N,\hat{T},\hat{\varphi})$, for every $v\in V$, there exists $x\in N_{v\preceq_{\hat{T}}}$ such that $d_{G[V_{\preceq_{\hat{T}}x}]}(x,v)\leq\Delta$. Let $Q$ be a shortest path in $G[V_{\preceq_{\hat{T}}x}]$ from $v$ to $x$. Note that $Q$ might contain vertices that are not in $G[V_{\preceq_{T}x}]$.
@@
Let $z\in N\cap Q$ such that $z$ has the highest order in $T$. By 3, every $y\in Q[v,z]$ satisfies $y\in V_{\preceq_{T}}z$. This implies $z\in N_{\preceq_{T}v}$ and $G[V_{\preceq_{T}z}](v,z)\leq w(Q[x,z])\leq w(Q)\leq\Delta$, giving the covering property. ∎
@@ 1 $\mathcal{R}\leftarrow\emptyset$
2 mark every vertex of $V$ uncovered
@@ 20      $\mathtt{round}\leftarrow\mathtt{round}+1$
Let $B_{1}$ and $B_{2}$ be the root bag of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, respectively. Assume w.l.o.g that $B_{1}$ is an ancestor of $B_{2}$. Let $X$ be the (only) leaf of $B_{1}$ that is the ancestor of $B_{2}$. By item 4 of 4, only the attachment of $X$ could possibly have a non-empty intersection $C_{\mathcal{T}_{2}}$. Suppose that there exists a vertex $v\in A(X)\cap C_{\mathcal{T}_{2}}$. Let $B$ be the child bag of $X$ that is an ancestor of $B_{2}$. Then $B\in\mathcal{T}^{\prime}$ and $B$ the center bag of a core $R$ created in round $i$. Observe that $v\in R$. When $R$ was created, all the uncovered vertices in $\mathcal{T}^{\prime}$ in $R$ were marked covered. This means $R\cap\mathtt{Uncov}(V[\mathcal{T}_{2}])=\emptyset$. Thus, $v\in R\cap A(Y)$ for some bag $Y\in\mathcal{T}^{\prime}$ in round $i+1$. However, by algorithm 2, $v$ will be removed from $A(Y)$ in round $i$ and hence will not be present in $A(Y)$ in round $i+1$, a contradiction. Thus, $C_{\mathcal{T}_{1}}\cap C_{\mathcal{T}_{2}}=\emptyset$. ∎

### v2

@@ Abstract
A $(\beta,\delta,\Delta)$-padded decomposition of an edge-weighted graph $G=(V,E,w)$ is a stochastic decomposition into clusters of diameter at most $\Delta$ such that for every vertex $v\in V$, the probability that $\mathbf{B}_{G}(v,\gamma\Delta)$ is entirely contained in the cluster containing $v$ is at least $e^{-\beta\gamma}$ for every $\gamma\in[0,\delta]$. Padded decompositions have been studied for decades and have found numerous applications, including metric embedding, multicommodity flow-cut gap, multicut, and zero extension problems, to name a few. In these applications, parameter $\beta$, called the padding parameter, is the most important parameter since it decides either the distortion or the approximation ratios. For general graphs with $n$ vertices, $\beta=\Theta(\log n)$.
@@ 1 Introduction
In this work, we study a stronger notion of stochastic decomposition, called padded composition. More formally, given a weighted graph $G=(V,E,w)$, a partition is $\Delta$-bounded if the diameter of every cluster is at most $\Delta$. A distribution $\mathcal{D}$ over partitions is called a $(\beta,\delta,\Delta)$-padded decomposition, if every partition is $\Delta$-bounded, and for every vertex $v\in V$ and $\gamma\in[0,\delta]$, we have:
$\Pr[\mathbf{B}_{G}(v,\gamma\Delta)\subseteq P(v)]\geq e^{-\beta\gamma}\qquad\text{where $P(v)$ is the cluster containing $v$.}$ (1)
@@
A key new technical insight to all of our aforementioned results is the notion of tree-ordered net (Definition 6). A tree order net is analogous to the notion of nets. Nets were used extensively in designing algorithms for metric spaces. More formally, a $\Delta$-net is a set of points $N$ such that every two net points are at a distance at least $\Delta$ (i.e $\min_{x,y\in N}d_{X}(x,y)\geq\Delta$), and every point has a net point at a distance at most $\Delta$ (i.e. $\max_{x\in V}\min_{y\in N}d_{X}(x,y)\leq\Delta$). Filtser [Fil19a] showed that if there is a $\Delta$-net such that every ball of radius $3\Delta$ contains at most $\tau$ net points, then the metric admits a $\left(O(\log\tau),\Omega(1),O(\Delta)\right)$-padded decomposition. This result implies that metrics of doubling dimension $d$ have padding parameter $O(d)$ since doubling metrics have sparse nets: $\tau=2^{O(d)}$. Unfortunately, graphs of small treewidth do not have sparse nets; this holds even in very simple graphs such as star graphs. Nonetheless, we show that small treewidth graphs possess a structure almost as good: a net that is sparse w.r.t. some partial order. We formalize this property via tree-ordered nets. As we will later show, a sparse tree-ordered net is enough to construct the padded decomposition scheme; see Section 3. We believe that the notion of a tree-ordered net is of independent interest.
A tree order $\preceq$ of a set $V$ is a partial order (i.e. transitive, reflexive, and anti-symmetric) associated with a rooted tree $T$ and a map $\varphi:V\rightarrow V(T)$ such that $u\preceq v$ iff $\varphi(v)$ is an ancestor of $\varphi(u)$ in $T$. Given a weighted graph $G=(V,E,w)$, a tree order $\preceq$ w.r.t. tree $T$ is a valid order of $G$ if for every edge $\{u,v\}\in E$, it holds that $u\preceq v$ or $v\preceq u$ or both (i.e. $v$ is an ancestor of $u$, or vice versa). A simple consequence of the validity is that every connected subset $C$ in $G$ must contain a maximum element w.r.t $\preceq$; see 11. For a vertex $v\in V$, and subset $S\subseteq V$ let $S_{v\preceq}=\{u\in S\mid v\preceq u\}$ be the ancestors of $v$ w.r.t. $T$ in $S$. Similarly, let $S_{\preceq x}=\{u\in S\mid u\preceq x\}$ be all the descendants of $x$ in $S$. See the illustration on the right.
@@ 2 Preliminaries
We consider connected undirected graphs $G=(V,E)$ with edge weights $w:E\to\mathbb{R}_{\geq 0}$. We say that vertices $v$ and $u$ are neighbors if $\{v,u\}\in E$. Let $d_{G}$ denote the shortest path metric in $G$. $\mathbf{B}_{G}(v,r)=\{u\in V\mid d_{G}(v,u)\leq r\}$ is the ball of radius $r$ around $v$. For a vertex $v\in V$ and a subset $A\subseteq V$, let $d_{G}(x,A):=\min_{a\in A}d_{G}(x,a)$, where $d_{G}(x,\emptyset)=\infty$. For a subset of vertices $A\subseteq V$, let $G[A]$ denote the induced graph on $A$, and let $G\setminus A:=G[V\setminus A]$.
@@
$\Pr[\mathbf{B}_{G}(z,\gamma\Delta)\subseteq P(z)]\geq e^{-\beta\gamma}~{}.$
@@ 3 Padded Decomposition and Sparse Cover
In this section, we prove three general lemmas constructing a padded decomposition and a sparse cover from a tree-ordered net. These lemmas, together with Lemma 8, imply Theorem 2, Theorem 3, and Theorem 5. We begin by proving Theorem 3, as its proof is simpler. We observe the following by the definition of a tree order.
Observation 11.
@@ 3.1 Proof of Theorem 3
$C_{i}=\mathbf{B}_{G[V_{\preceq x_{i}}]}(x_{i},\alpha\Delta)~{},$
@@
Finally, we show that for every vertex $v$, $\mathbf{B}_{G}(v,\frac{\Delta}{\beta})$ is fully contained in some cluster in $\mathcal{C}$, for $\beta=\frac{2}{\alpha-1}$. Let $B=\mathbf{B}_{G}(v,\frac{\Delta}{\beta})$. Let $v_{B}\in B$ be the closest vertex to the root w.r.t. $T$. 11 implies that for every vertex $u\in B$, $u\preceq v_{B}$. By the triangle inequality, for every $u\in B$ it holds that:
@@ 10             Add $C_{x}$ to $\mathcal{P}_{i}$
$d_{G}(x,x_{i})\leq d_{G[V_{\preceq x}]}(x,y_{i})+d_{G[V_{\preceq x_{i}}]}(y_{i},x_{i})\leq\alpha\Delta~{}.$
@@
$\leq\Delta+\frac{2\Delta}{\beta}=\left(1+\frac{2\cdot(\alpha-2)}{4}\right)\cdot\Delta=\frac{\alpha}{2}\cdot\Delta~{},$
@@ 3.3 Proof of Theorem 2
$C_{i}=\mathbf{B}_{G[V_{\preceq x_{i}}]}(x,R_{i})\setminus\cup_{j<i}C_{j}~{}.$
@@
Let $v_{B}\in B$ be the vertex closest to the root of $T$ (w.r.t distances in $T$). By 11, for every vertex $u\in B$, it holds that $u\preceq v_{B}$. Let $x_{B}\in N$ be the center that is an ancestor of $v_{B}$ and minimizes $\mathbf{B}_{G[V_{\preceq x}]}(x,v_{B})$. Note that $v_{B}$ will join the cluster of $x_{B}$, if it did not join any other cluster. It follows that no descendant of $x_{B}$ can be the center of the first cluster having a non-trivial intersection with $B$. This implies that for every center $x_{i}\notin N_{x_{B}\preceq}$, $\Pr[\mathcal{F}_{i}]=0$.
@@
We assume that by the round $i$, no vertex in $B$ is clustered, and that $d_{G[V_{\preceq x_{i}}]}(x_{i},B)\leq\beta\Delta$, as otherwise $\Pr[\mathcal{C}_{i}]=\Pr[\mathcal{F}_{i}]=0$ and the proof follows. Let $\rho$ be the minimal value of $\delta_{i}$ such that if $\delta_{i}\geq\rho$, some vertex of $B$ will join $C_{i}$. Formally $\rho=\frac{1}{\Delta}\cdot d_{G[V_{\preceq x_{i}}]}(x_{i},B)$. By our assumption, $\rho\leq\beta$. Set $\tilde{\rho}=\max\{\rho,1\}$. We have:
@@
$d_{G[V_{\preceq x_{i}}]}(u,x_{i})\leq d_{G[V_{\preceq x_{i}}]}(v_{i},x_{i})+2\gamma\Delta=\left(\rho+2\gamma\right)\cdot\Delta~{}.$
@@
$=\left(1-e^{-2\gamma\cdot\lambda}\right)\cdot\left(\Pr\left[\mathcal{F}_{i}\right]+\frac{1}{e^{(\beta-1)\cdot\lambda}-1}\right)~{}.$
@@
$\leq\left(1-e^{-2\gamma\cdot\lambda}\right)\cdot\left(1+e^{-2\gamma\cdot\lambda}\right)=1-e^{-4\gamma\cdot\lambda}~{},$
@@
$e^{-2\gamma\lambda}=\frac{e^{-2\gamma\lambda}\left(e^{(\beta-1)\cdot\lambda}-1\right)}{e^{(\beta-1)\cdot\lambda}-1}\geq\frac{e^{-2\gamma\lambda}\cdot e^{(\beta-1)\cdot\lambda}\cdot\frac{1}{2}}{e^{(\beta-1)\cdot\lambda}-1}\overset{(*)}{\geq}\frac{\frac{1}{2}\cdot e^{\frac{\alpha-1}{4}\cdot\lambda}}{e^{(\beta-1)\cdot\lambda}-1}=\frac{\tau}{e^{(\beta-1)\cdot\lambda}-1}~{}.$
@@ 4 Tree-Ordered Nets for Graphs of Bounded Tree-partition Width
The converse of 18 might not be true, specifically in the case when $u$ and $v$ are mapped to the same vertex in $\hat{T}$ by $\hat{\varphi}$. By 18, for any vertex $x\in V$, $V_{\preceq_{T}x}\subseteq V_{\preceq_{\hat{T}}x}$. Thus, for any $v$ and $x$ such that $v\preceq_{T}x$, if $d_{G}[V_{\preceq_{T}x}](v,x)\leq 3\Delta$, then $d_{G}[V_{\preceq_{\hat{T}}x}](v,x)\leq 3\Delta$. Therefore, $N^{3\Delta}_{v\preceq_{T}}\subseteq N^{3\Delta}_{v\preceq_{\hat{T}}}$ and hence $|N^{3\Delta}_{v\preceq_{T}}|\leq|N^{3\Delta}_{v\preceq_{\hat{T}}}|=\mathrm{poly}({\rm tp})$, giving the packing property.
Next, we show the covering property. By the covering property of $(N,\hat{T},\hat{\varphi})$, for every $v\in V$, there exists $x\in N_{v\preceq_{\hat{T}}}$ such that $d_{G[V_{\preceq_{\hat{T}}x}]}(x,v)\leq\Delta$.
Let $Q$ be a shortest path in $G[V_{\preceq_{\hat{T}}x}]$ from $v$ to $x$. Note that $Q$ might contain vertices that are not in $G[V_{\preceq_{T}x}]$.
@@
Let $z\in N\cap Q$ such that $z$ has the highest order in $T$. By 19, every $y\in Q[v,z]$ satisfies $y\in V_{\preceq_{T}}z$. This implies $z\in N_{\preceq_{T}v}$ and $G[V_{\preceq_{T}z}](v,z)\leq w(Q[x,z])\leq w(Q)\leq\Delta$, giving the covering property. ∎
@@ 1 $\mathcal{R}\leftarrow\emptyset$
2
3mark every vertex of $V$ uncovered
@@ 21      $\mathtt{round}\leftarrow\mathtt{round}+1$
Let $B_{1}$ and $B_{2}$ be the root bag of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, respectively. Assume w.l.o.g that $B_{1}$ is an ancestor of $B_{2}$. Let $X$ be the (only) leaf of $B_{1}$ that is the ancestor of $B_{2}$. By item 4 of 20, only the attachment of $X$ could possibly have a non-empty intersection $C_{\mathcal{T}_{2}}$. Suppose that there exists a vertex $v\in A(X)\cap C_{\mathcal{T}_{2}}$. Let $B$ be the child bag of $X$ that is an ancestor of $B_{2}$. Then $B\in\mathcal{T}^{\prime}$ and $B$ the center bag of a core $R$ created in round $i$. Observe that $v\in R$. When $R$ was created, all the uncovered vertices in $\mathcal{T}^{\prime}$ in $R$ were marked covered. This means $R\cap\mathtt{Uncov}(V[\mathcal{T}_{2}])=\emptyset$. Thus, $v\in R\cap A(Y)$ for some bag $Y\in\mathcal{T}^{\prime}$ in round $i+1$. However, by algorithm 2, $v$ will be removed from $A(Y)$ in round $i$ and hence will not be present in $A(Y)$ in round $i+1$, a contradiction. Thus, $C_{\mathcal{T}_{1}}\cap C_{\mathcal{T}_{2}}=\emptyset$. ∎